The ideal height (in inches) of a man is related to his weight(in pounds) by the formula:
H=W/5+190/5
Solve for W and calculate how much a man of height 62 inches should weigh.

Answer by checkley71(8403)

H=W/5+190/5
62=W/5+190/5
W/5=62-190/5
W/5=(310-190)/5
W=120